### v3

# Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games

Abstract
We present Questum, a novel framework for Large Language Model (LLM)-based agents in Murder Mystery Games (MMGs). MMGs pose unique challenges, including undefined state spaces, absent intermediate rewards, and the need for strategic interaction in a continuous language domain. Questum addresses these complexities through a sensor-based representation of agent states, a question-targeting mechanism guided by information gain, and a pruning strategy to refine suspect lists and enhance decision-making efficiency. To enable systematic evaluation, we propose WellPlay, a dataset comprising 1,482 inferential questions across 12 games, categorised into objectives, reasoning, and relationships. Experiments demonstrate Questum’s capacity to achieve superior performance in reasoning accuracy and efficiency compared to existing approaches, while also significantly improving the quality of agent-human interactions in MMGs. This study advances the development of reasoning agents for complex social and interactive scenarios.
Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games
Qinglin Zhu1∗, Runcong Zhao1∗, Jinhua Du2, Lin Gui1, Yulan He1,3 1King’s College London, 2Huawei London Research Centre, 3The Alan Turing Institute {qinglin.1.zhu, runcong.zhao}@kcl.ac.uk{jinhua.du}@huawei.com {lin.1.gui,yulan.he}@kcl.ac.uk
**footnotetext: Equal contribution.
1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023). Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite these advances, building agents for Murder Mystery Games (MMGs) that intentionally interact with humans in dynamic environments remains challenging.
Figure 1: The murder mystery game involves players taking on roles, questioning suspects, and using deduction to identify the killer, all while pursuing their character-specific objectives.
As shown in Figure 1, MMGs are strategic games that involve 4-12 players assuming character roles with specific scripts and objectives. Gameplay unfolds through language-based negotiation and tactical coordination over $n$ rounds, with players limited to $m$ questions per round. The ultimate goal is to uncover the truth collaboratively while advancing individual objectives. At the end of the game, players vote on the suspected killer and evaluate their objective completion.
MMGs are script-based games in which each script can only be played once. Unlike traditional games such as Go, which have fixed settings and well-defined action spaces for searching, or Werewolves (e.g., voting results) and Avalon (e.g., mission success), which have intermediate rewards, MMGs lack both well-defined searching spaces and intermediate rewards in decision making.
To address these challenges, we first propose using a discrete space, which is called "sensor" to model each character to character states. Since it can be captured or updated by prompting the natural language description, we are able to explore this search space using generative Question-Answering (what to ask) between agents in real time as a regular planning problem (Karaman and Frazzoli, 2011; Gammell et al., 2020). For the issue of intermediate reward in searching, we design an Information Gain (IG) based heuristic objective which aims to balance existing searching results and the expected reward for further questions. After each action, the agent evaluates whether the current information can update the character states and narrow down the range of suspects. Meanwhile, our method focuses on extracting highly suspicious characters, creating a question-target selection framework guided by IG (who to ask). Compared to previous work, this framework optimises the information collection path, significantly improving system efficiency.
Furthermore, current evaluation methods either focus solely on the final win/loss outcome or rely heavily on manual evaluation, which significantly limits the scope of analysis due to the high-cost (Xu et al., 2023a; Wang et al., 2023; Wu et al., 2024). For instance, Wu et al. (2024) manually designed only 56 inferential questions for 4 selected MMGs. To address the evaluation problem, we constructed a dataset, WellPlay, containing 1,482 inferential questions across 12 MMGs. These questions are categorised into three types: objective, reasoning, and relation. Including the win rate metric used in previous works, this provides a more diverse evaluation framework. We also recruited human players to interact with the agent. The results revealed that previous agents often overfit to agent-vs-agent interaction environments, exhibiting verbose and repetitive dialogues that detracted from the player experience. By integrating our goal-oriented pruner, these issues were significantly mitigated, resulting in more human-like agent behaviour and greatly enhancing player satisfaction.
In summary, we have made the following contributions: (1) We propose Questum, a framework using detective sensors to define and optimise the search in continuous language space. (2) We introduce an IG-based reward to address the lack of intermediate rewards, improving agent behaviour and efficiency. (3) We construct WellPlay, a dataset with 1,482 inferential questions across diverse categories, offering a more robust evaluation framework. (4) By integrating a goal-oriented pruner, we enhance agent-human interaction, reducing verbosity and improving player experience.  [^1] Our code and dataset are available at https://github.com/alickzhu/Questum.
2 Questum
In this section, we describe the MMG settings and provide an overview of how Questum works.
2.1 Problem Setting
In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored to such scenarios. This framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ and a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{a}$ and $N_{v}$ denote the numbers of playable characters and victims, respectively. Each agent $a_{i}$ is assigned to a character and initialised with the following:
(1) Role $\mathbf{r}_{i}=\{r_{ik}\}_{k=1}^{N_{v}}$: Whether or not they are the murderer of each victim;
(2) Role background script $C_{i}$: Crafted from the unique viewpoint of $a_{i}$;
(3) Objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{o_{i}}}$: A set of $N_{o_{i}}$ goals for agent $a_{i}$.
Game Rules are also provided to Agents as essential information. The detailed rules and procedures are provided in Appendix A.
2.2 Questum  Planning Strategy
Figure 2: Search and Approximate. Questum  generates questions based on character states, selecting agents to question based on past observations of critical information and the likelihood of uncovering more. The goal is to minimise the suspect list.
Questum approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 2, this framework is fundamentally composed of two key components: (1) Search based on sensors: Questum  searches for the murderer based on their proximity to the ideal murderer in the language space. (2) Approximation with a pruner: Questum  focuses on a subset of characters that are highly suspicious of being a murderer, and decides whom to query based on this suspect list.
In MMGs, the suspect status is $\mathbf{s_{i}}=\left[s_{ijk}\right]_{j=1,k=1,j\neq i}^{N_{a},N_{v}}$, where $s_{ijk}\in[0,1]$. Here, $s_{ijk}$ encodes agent $a_{i}$’s beliefs about player $a_{j}$’s likelihood of being victim $v_{k}$’s murderer. In our setting, $s_{ijk}$ is set to be discrete ($\in\{0,1\}$), representing suspect and non-suspect. While our list of suspects is finite, our actual search space $L\in\mathbb{R}^{d}$ (the language space) is continuous and high-dimensional, where $d$ is potentially very large. It encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets.
Search based on sensors
We map each suspect to a point in the state space $L$. Our search process operates continuously in the state space $L$, where character attributes such as emotions, motivations, and opportunities are represented as vectors $u_{i}=(e_{i},m_{i},p_{i},...)$ in this multi-dimensional space. Character attributes are chosen with domain knowledge, drawing on insights from sociology (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010). We use sensors to sample the values of these attributes for two main purposes:
1) Estimating the distance between a given suspect and the hypothetical murderer. The hypothetical murderer is assumed to be a point in the search space. At the beginning of the game, when participants lack additional information, the hypothetical murderer refers to an idealised profile that meets specific criteria across all sensor dimensions: someone with motivation, negative emotions, and the opportunity to commit the crime. As the game progresses and more available information, agents update the criteria defining the hypothetical murderer. For instance, if it is suddenly revealed that the victim’s death was accidental, the murderer may no longer be expected to have a motive.
2) Exploration by generating questions. It is modelled as $q=g(u_{i},X)$, where $q$ is a newly generated question by language model $g$. In the inputs, $X$ is the relevant information retrieved from the context by Retrieval-Augmented Generation (RAG) and $u_{i}$ is the state representation of suspect $i$.
Approximation with a pruner
Following the questioning step, agents receive responses from the questioned agent as well as conversational exchanges among other agents. Each agent then updates its understanding of interpersonal relationships, thereby updating character states. In human gameplay, players often form judgments regarding the scenario and focus on a subset of highly suspicious individuals. This strategic refinement process is crucial for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Inspired by the human thinking process, we also enable agents to consider the suspect list.
In an MMG with $N_{a}$ players, to find out who is the murderer $Y$ of a given victim $v_{k}$, we define the entropy to be $H=-ln(\frac{1}{n})$, where $n$ is the number of suspects in the agent’s suspect list. Based on the size of the suspect list for each round of the game, the agent can calculate the corresponding entropy. The target is to maximise this entropy. The initial entropy is $H_{0}=-ln(\frac{1}{N_{a}-1})$, where everyone except the agent itself, who aims to find the murderers, are all suspects, representing the uncertainty about who the murderer is at the start of the game. For each character, we have its historical IG being $IG_{i}=H_{i-1}-H_{i}$ for round $i$, which represents the reduction in uncertainty about the murderer’s identity after receiving the sensor reading and the corresponding answer. $IG_{1:i}=[IG_{1},IG_{2},...,IG_{i}]$. The weighted historical IG is:
$IG_{weighted}[i]=\frac{\sum_{k}w_{k}\times IG_{k}}{\sum_{k}w_{k}}$ (1)
where $w_{k}$ are weights, set to $e^{-(i-k)}$.
After the search step, the agent acquires more information. However, accessing the conditional probability to update the conditional entropy $H_{i+1}=H(Y|u_{i},X)$ might be challenging in closed-source platforms like ChatGPT, and open-source projects such as Llama could introduce strong bias as they usually focus on one suspect Xu et al. (2023b). Hence, we propose a heuristic estimator based on prompting. First, we ask the LLMs to estimate the expected IG. We then convert the discrete value from the prompt to a numerical value:
$E[IG_{i+1}]=\begin{cases}1,&\text{if prompt returns ``high''}\\ 0,&\text{if prompt returns ``medium''}\\ -1,&\text{if prompt returns ``low''}\end{cases}$ (2)
Thus, even closed-source platforms are able to estimate conditional entropy, and the bias in open-source models is reduced by smoothing in the form of multiple-choice questions. We then combine historical and expected gain to determine the highest-scoring character to ask:
$Score_{i}=\beta IG_{weighted}[i]+(1-\beta)E[IG_{i+1}]$ (3)
In addition, we also implement an $\epsilon$-greedy strategy to choose a random character with probability $\epsilon$, and the highest-scoring character with probability $1-\epsilon$ to balance exploration and exploitation. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. Thus, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and increasing the likelihood of identifying the perpetrator within the designated number of discussion rounds. The detailed procedure is in Algorithm 1.
Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, Suspicious States $s$, Max round $n$.
Output: Evaluation of Results
1 current_round = 0
2 while current_round $\leq$ $n$ do
       // Search based on sensors
3       for i = 1 to $N_{a}$ do
4             for k = 1 to $N_{v}$ do
5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
6                   for $a_{j}$ in $suspect\_list_{ik}$ do
7                         question = Action_Generation($a_{i},a_{j},v_{k}$)
8                         answer = Reply($a_{i}$)
9
10
11
      // Approximation with a pruner
12       for i = 1 to $N_{a}$ do
13             for k = 1 to $N_{v}$ do
14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{k}$)
15
16
17      current_round += 1
Algorithm 1 Questum Framework
3 The WellPlay Dataset
MMG #Agents #Victims #token(CN) #token(EN) Question
avg overall avg overall Objective Reasoning Relations overall
Death Wears White 9 1 3,191 28,716 1,742 15,681 10 102 72 184
Ghost Revenge 7 3 5,488 38,415 3,960 27,723 19 152 69 240
Danshui Villa 7 2 5,111 35,779 3,339 23,370 12 128 63 203
Unfinished Love 7 2 2,501 17,507 1,652 11,562 12 61 72 145
Cruise Incident 5 1 1,263 6,313 808 4,040 4 24 30 58
Sin 4 1 2,121 8,485 1,378 5,512 3 20 21 44
Deadly Fountain 4 1 1,852 7,410 1,194 4,775 3 21 12 36
Unbelievable Incident 5 1 3,182 15,912 2,012 10,062 4 24 15 43
Desperate Sunshine 4 1 3,370 13,481 2,219 8,874 3 18 36 57
Riverside Inn 4 1 1,910 7,638 1,257 5,028 3 18 18 39
Solitary Boat Firefly 6 4 8,894 53,362 6,874 41,244 20 109 69 198
Manna 6 3 9,028 54,169 6,492 38,954 24 123 88 235
Avg 5.67 1.75 3,993 23,932 2,744 16,402 9.75 66.67 47.08 125.50
Sum 68 21 47,911 287,187 32,927 196,825 117 800 565 1482
Table 1: Dataset Statistics. Agents is the count of players, Victims is the number of victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types. The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
Dataset
We built an evaluation dataset, WellPlay, derived from background narratives created for MMGs (Zhao et al., 2024c). To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers. We have recruited four annotators to label inferential questions on:
(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles the wallet, for each character in the game.
(2) Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, or disappearance); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, or manslaughter).
(3) Relations. This includes interpersonal relationships between victims and others, as well as relationships among suspects, with labels adapted from the Conan dataset (Zhao et al., 2024c).
WellPlay encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims (see Table 1). More details of the dataset and annotating schema are in Appendix B.
4 Experiments
Script Evaluation Werewolf O-CoT ThinkThrice Questum
Death Wears White (9 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$033_{\pm.047}.$$067_{\pm.047}.$$067_{\pm.047}.$$033_{\pm.047}.$
Reasoning .$258_{\pm.024}.$$324_{\pm.014}.$$327_{\pm.012}.$$356_{\pm.030}.$
Relations .$421_{\pm.024}.$$458_{\pm.041}.$$384_{\pm.026}.$$435_{\pm.017}.$
Ghost Revenge (7 players, 3 victims) Win Rate .$222_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$333_{\pm.000}.$
Objective .$193_{\pm.066}.$$158_{\pm.043}.$$193_{\pm.025}.$$333_{\pm.066}.$
Reasoning .$307_{\pm.008}.$$322_{\pm.019}.$$377_{\pm.022}.$$353_{\pm.016}.$
Relations .$314_{\pm.018}.$$295_{\pm.018}.$$353_{\pm.025}.$$353_{\pm.018}.$
Danshui Villa (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$083_{\pm.000}.$$111_{\pm.039}.$$194_{\pm.039}.$$111_{\pm.079}.$
Reasoning .$286_{\pm.027}.$$286_{\pm.019}.$$310_{\pm.004}.$$286_{\pm.004}.$
Relations .$312_{\pm.027}.$$365_{\pm.045}.$$259_{\pm.015}.$$296_{\pm.075}.$
Unfinished Love (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$500_{\pm.000}.$
Objective .$083_{\pm.118}.$$028_{\pm.039}.$$000_{\pm.000}.$$333_{\pm.068}.$
Reasoning .$443_{\pm.023}.$$426_{\pm.035}.$$481_{\pm.039}.$$536_{\pm.028}.$
Relations .$472_{\pm.034}.$$519_{\pm.036}.$$514_{\pm.020}.$$560_{\pm.007}.$
Cruise Incident (5 players, 1 victim) Win Rate .$667_{\pm.471}.$ $1.000_{\pm.000}$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective .$417_{\pm.118}.$$500_{\pm.000}.$$583_{\pm.312}.$$667_{\pm.236}.$
Reasoning .$458_{\pm.068}.$$444_{\pm.086}.$$458_{\pm.059}.$$528_{\pm.052}.$
Relations .$367_{\pm.082}.$$422_{\pm.042}.$$411_{\pm.016}.$$422_{\pm.016}.$
Sin (4 players, 1 victim) Win Rate .$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$
Objective .$333_{\pm.272}.$$000_{\pm.000}.$$000_{\pm.000}.$$444_{\pm.314}.$
Reasoning .$650_{\pm.108}.$$467_{\pm.024}.$$533_{\pm.047}.$$550_{\pm.041}.$
Relations .$333_{\pm.067}.$$571_{\pm.067}.$$413_{\pm.090}.$$492_{\pm.022}.$
Deadly Fountain (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Reasoning .$381_{\pm.000}.$$444_{\pm.098}.$$508_{\pm.022}.$$587_{\pm.022}.$
Relations .$250_{\pm.068}.$$194_{\pm.039}.$$222_{\pm.039}.$$333_{\pm.068}.$
Unbelievable Incident (5 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$083_{\pm.118}.$$000_{\pm.000}.$$000_{\pm.000}.$$083_{\pm.118}.$
Reasoning .$431_{\pm.052}.$$278_{\pm.071}.$$194_{\pm.071}.$$472_{\pm.020}.$
Relations .$622_{\pm.083}.$$733_{\pm.109}.$$289_{\pm.126}.$$533_{\pm.054}.$
Desperate Sunshine (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective .$333_{\pm.000}.$$111_{\pm.157}.$$556_{\pm.157}.$$778_{\pm.157}.$
Reasoning .$537_{\pm.026}.$$630_{\pm.094}.$$759_{\pm.052}.$$741_{\pm.069}.$
Relations .$491_{\pm.094}.$$574_{\pm.035}.$$491_{\pm.047}.$$556_{\pm.068}.$
Riverside Inn (4 players, 1 victim) Win Rate .$083_{\pm.118}.$$000_{\pm.000}.$$083_{\pm.118}.$$250_{\pm.204}.$
Objective .$117_{\pm.024}.$$117_{\pm.062}.$$117_{\pm.094}.$$383_{\pm.062}.$
Reasoning .$248_{\pm.013}.$$312_{\pm.022}.$$281_{\pm.011}.$$373_{\pm.017}.$
Relations .$531_{\pm.014}.$$541_{\pm.038}.$$483_{\pm.030}.$$589_{\pm.071}.$
Solitary Boat Firefly (6 players, 4 victims) Win Rate .$222_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective .$250_{\pm.090}.$$181_{\pm.039}.$$181_{\pm.039}.$$250_{\pm.034}.$
Reasoning .$409_{\pm.004}.$$369_{\pm.044}.$$453_{\pm.008}.$$539_{\pm.027}.$
Relations .$473_{\pm.028}.$$568_{\pm.048}.$$439_{\pm.033}.$$542_{\pm.027}.$
Manna (6 players, 3 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$333_{\pm.471}.$
Objective .$111_{\pm.157}.$$000_{\pm.000}.$$556_{\pm.157}.$$444_{\pm.157}.$
Reasoning .$463_{\pm.052}.$$426_{\pm.026}.$$593_{\pm.026}.$$648_{\pm.069}.$
Relations .$444_{\pm.045}.$$407_{\pm.069}.$$333_{\pm.045}.$$444_{\pm.045}.$
Overall Win Rate .$127_{\pm.059}.$$063_{\pm.022}.$$111_{\pm.045}.$$222^{*}_{\pm.081}.$
Objective .$160_{\pm.020}.$$123_{\pm.008}.$$162_{\pm.024}.$$288^{*}_{\pm.021}.$
Reasoning .$343_{\pm.006}.$$349_{\pm.012}.$$384_{\pm.002}.$$423_{\pm.009}.$
Relations .$425_{\pm.008}.$$473_{\pm.010}.$$405_{\pm.006}.$$471_{\pm.009}.$
Overall .$324_{\pm.004}.$$329_{\pm.008}.$$347_{\pm.006}.$$407^{*}_{\pm.006}.$
Table 2: Results of Agent-vs-Agent Evaluation. An “*” indicates statistical significance under the two-sample t-test with a significance level of $\alpha=0.05$, comparing it with the second-best model. More detailed data is provided in Table 8.
4.1 Experimental Setup
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models , using gpt-35-turbo-16k 0613 and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and standard deviation. In this paper, we focus on the strategy for the good camp. The murderer agent’s questioning strategy mirrors that of other agents, with an added prompt instructing it to act as if it is not the murderer while questioning others.
4.2 Evaluation Metrics
For evaluation, we adopt the following metrics:
Win Rate:
Every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) is eliminated. Murderers always vote for others instead of themselves. If the true murderer is voted out, the game is considered a victory for the identifying players.
Question Accuracy:
The accuracy of agents’ answers to three types of questions—Objective, Reasoning, and Relations—and Overall, where Overall represents the combined accuracy across all question types to provide a comprehensive evaluation of the agents’ performance. This metric is used in both Agent-vs-Agent and Human-vs-Agent evaluations to measure agents’ ability to answer inferential questions correctly.
4.3 Baselines
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games. Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games.
(1) Werewolf (Xu et al., 2023a). Werewolf is another multi-agent game, where players identify werewolves through group discussion. Questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. We adapt its pre-set game instructions and role-specific information to MMGs settings. (2) Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b). Agents think, reflect, and choose who and what to ask based on their objectives. We use the framework from previous works, only replacing the agents’ objectives with those set in MMGs. (3) ThinkThrice (Wu et al., 2024). Designed for MMGs, agents craft questions from retrieved memory and the current scenario. (4) Personal Perspective (PP). For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script. (5) Omniscient Perspective (OP). Agents do not actively participate but make their final decisions based on all agents’ scripts.
To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and Questum). The implementation details, model comparison, information on sensors, method to calculate the Overall, and prompts are provided in Appendix C.
4.4 Agent-vs-Agent Evaluation
Figure 3: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. The Personal Perspective (PP) is designed to represent the starting point for searching. The Omniscient Perspective (OP) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint.
In this section, we evaluate agent performance in an Agent-vs-Agent setting to compare their decision-making processes against competitive baselines. Figure 3 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table 2). Questum  shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Notably, Questum  significantly outperforms others in Objective Questions. The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies. It represents the ideal “search” endpoint. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model. This showcases the “approximate” ability of Questum  in refining information and dynamically narrowing down the search domain to achieve the target objective. Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
The results demonstrate that Questum  outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability. The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals. This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
We also evaluate the performance of our framework on English scripts and open-source LLMs in Appendix D. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
4.5 Agent-vs-Human Evaluation
In the Agent-vs-Human Evaluation, we identified significant limitations of traditional evaluation methods used in agent for games, which were primarily designed for Agent-vs-Agent Evaluation and focused solely on achieving victory. While effective in assessing agents’ strategic and reasoning abilities, these methods fall short when considering human players as the end-users of the system. For instance, previous approaches often resulted in repetitive and meaningless conversations, which human players found unengaging and boring. To illustrate these limitations, we provide specific examples in Appendix E, drawn from dialogues that occurred during games involving human participants.
To address this, we extended our evaluation by incorporating a player-centric perspective. In addition to the performance metrics used in Agent-vs-Agent Evaluation, we distributed a survey to human participants to capture their gameplay experience and satisfaction levels. This allowed us to assess Questum’s suitability not only as a strategic agent but also as a companion for human players in interactive settings. The metrics include:
Story Advancement: Measures the agent’s effectiveness in gathering relevant information and advancing the game’s plot in a meaningful way. Higher scores indicate that the interactions were productive and kept the game engaging.
Question Quality: Evaluates the relevance and depth of the questions posed by the agent. This metric ensures that the questions contribute to the narrative and help uncover critical information, avoiding redundancy or triviality.
Response Quality: Assesses the accuracy and informativeness of the agent’s responses. Higher scores indicate that the agent provided clear and valuable information to the human players.
Response Speed: Represents how quickly the agent interacts, ensuring timely response for a smooth and engaging gameplay experience.
Role Immersion: Assesses how well the agent embodies its character’s personality, emotions, and background to create a believable experience.
Evaluation Werewolf O-CoT ThinkThrice Questum
Win Rate .$111_{\pm.157}.$$333_{\pm.132}.$$333_{\pm.048}.$$667_{\pm.085}.$
Objective .$222_{\pm.091}.$$333_{\pm.045}.$$296_{\pm.052}.$$556_{\pm.091}.$
Reasoning .$678_{\pm.028}.$$644_{\pm.028}.$$684_{\pm.008}.$$723_{\pm.035}.$
Relations .$745_{\pm.028}.$$797_{\pm.040}.$$745_{\pm.028}.$$843_{\pm.042}.$
Overall .$608_{\pm.006}.$$619_{\pm.023}.$$625_{\pm.009}.$$717_{\pm.038}.$
Table 3: Performance comparison of different algorithms against human players.
Figure 4: Results of Agent-vs-Human Evaluation using Human-Centric metrics. The detailed data can be found in Table 6.
All scores are based on discrete 1-5 scales, where higher scores are better. The specific scoring criteria can be found in Table 7. Details about the setups (e.g., recruitment of players, gameplay conditions, and tasks) can be found in Appendix B.5.
As shown in Table 3 (results against human players) and Figure 4 (Human-Centric metrics), Questum  not only surpassed the baseline models in strategic and reasoning-oriented evaluation but also demonstrated superior performance in human-centric dimensions. These results indicate that Questum  effectively balances tactical prowess with narrative engagement, delivering more satisfying gameplay experiences for human participants.
4.6 Efficiency and Cost Analysis
As shown in Figure 5, we delve into the efficiency and cost analysis across various methodologies for agent interaction in MMGs. The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense. [^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Figure 5: Comparison of time (hours) and costs (\$) for calling OpenAI API across multi-agent algorithms in MMG settings.
We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps. In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives. Questum  stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, Questum  minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with Questum  consistently registering lower costs in comparison to its counterparts during the Gameplay phase. The trend in time efficiency mirrors that of the monetary cost, with similar patterns observed across the methodologies.
4.7 Ablation Studies
We conducted comprehensive ablation studies to evaluate the effectiveness of the key components of our approach. Removing the Pruner or Sensor modules adversely affected both performance and cost reduction, confirming their utility; results are provided in Table 12. Additionally, we experimented with different sensor configurations and numbers of rounds/questions, results are shown in Figure 7. Our analyses revealed that a carefully chosen subset of three sensors strikes a balance between performance gains and computational costs and that around three rounds of interaction with one question per round is near-optimal. These experimental setups, alongside full performance comparisons and in-depth evaluations, are presented in the Appendix D.3.
5 Related Works
Multi-Agent Interaction
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. Compared to classic methods (Wang and Shen, 2024), LLMs based methods can perform inference across a wider variety of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, the inherit biases that can potentially limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024; Wang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
Optimisation for Complicated Tasks
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. Therefore, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approaches was also adapted to search for optimal tools (Du et al., 2024), and interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024). We were also inspired by stochastic search methods for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
6 Conclusion
This study introduces Questum, a framework that tackles the challenges of LLM-driven agents in MMGs through sensor-based modeling, information gain-driven questioning, and suspect pruning. With the WellPlay dataset enabling robust evaluation, Questum  demonstrates superior reasoning, efficiency, and human-like interaction quality. These advancements pave the way for more effective and engaging AI systems in dynamic multi-agent scenarios.
Limitation
Annotation Scope
Murder Mystery Games are built on intricately written human-authored stories, featuring dense plots, complex character relationships, and unexpected twists to enhance gameplay. These characteristics make annotation particularly challenging, explaining the scarcity of labeled datasets for MMGs. Existing resources include only 56 inferential questions across 4 selected MMGs.
While this dataset is still insufficient for training due to the task’s inherent complexity—and the impracticality of creating a large dataset for such scenarios—we believe it adequately demonstrates the effectiveness of our methodology. Future research could explore synthetic data generation with enhanced quality control to further expand the dataset.
Generalisation Across Domains
While the proposed framework demonstrates strong performance for MMGs, its generalisability to other multi-agent or language-driven tasks has not been fully explored. Applying this approach to domains with different interaction dynamics or less structured narratives may require substantial modifications.
Future research could extend the framework to other multi-agent scenarios, such as collaborative storytelling, negotiation games, or real-world applications like customer service. Domain adaptation techniques could help refine the framework for broader use cases.
Resource Dependence
The framework depends on powerful LLMs for sensor-based reasoning and question generation, which can be resource-intensive. Scaling this approach to larger or real-time systems may require additional computational and memory resources, limiting accessibility for smaller-scale applications.
Efforts to optimise computational efficiency, such as caching mechanisms, or lightweight models for real-time applications, could make the framework more accessible and scalable.
Ethics Statement
Please note that MMGs and our WellPlay dataset, created to assess agents’ behaviors in MMGs, may include descriptions of violent events, actions, or characters. This content is included solely for academic, research, and narrative analysis purposes. It is not meant to glorify or trivialise violence in any form. We have informed annotators about the potential exposure to violent content. Users or researchers intending to use this dataset should be aware of this potential exposure and are advised to engage with the dataset in a professional and responsible manner. This dataset is unsuitable for minors and those sensitive to such content.
Acknowledgments
This work was supported in part by the UK Engineering and Physical Sciences Research Council through a New Horizons grant (EP/X019063/1) and a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2).
Appendix A MMGs Rules and Procedure
A.1 Detailed Rules
Rule 1:
The total number of players participating in the game depends on the script. There may be one or more players who are the murderer(s), while the rest are civilians.
Rule 2:
The goal of the game is for civilian players to collaborate and face a meticulously planned murder case together, collecting evidence and reasoning to identify the real murderer among the suspects, all the while ensuring they are not mistaken for the murderer; murderer players must concoct lies to hide their identity and avoid detection, while also achieving other objectives in the game.
Rule 3:
Throughout the game, only murderer players are allowed to lie. To conceal their identity, murderers may choose to frame others to absolve themselves of guilt; non-murderer players (civilians) must answer questions from other players and the host honestly and provide as much information as they know about the case to help uncover the truth.
Rule 4:
At the start of the game, each player receives their character script from the host, which contains information about their role and identity.
Rule 5:
Other players cannot see the content of each player’s character script, so players must and can only collect information about other players through interaction after the game starts.
Rule 6:
In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with over 50% of the votes is the murderer, the civilian players win. Otherwise, the murderer players win.
A.2 Procedure
Stage 1: Distribution of Character Scripts
The host distributes character scripts to each player. These scripts contain the player’s name, role (murderer or civilian), and a brief character backstory.
Stage 2: Self-Introduction Session
Players introduce their characters to the group, laying the groundwork for the game’s interactions.
Stage 3: Rounds of Open Questioning
The game progresses through three rounds of open questioning. Players take turns to ask and answer questions, aiming to gather information about others.
Stage 4: Voting
In this stage, players vote anonymously to determine their suspicions regarding the identity of the murderer. Each player has one vote.
Stage 5: Outcome Reveal
The game concludes with the announcement of the voting results, revealing whether the civilian players successfully identified the murderer or not.
A.3 Example: Solitary Boat Firefly Script
As an illustrative example, we examine the Solitary Boat Firefly murder mystery script, involving six players: [ “Tian Chou", “Zhou Lianyi", “Xi Yan", “Yu Sunian", Yannan", and “Zhou Chitong" ], along with four victims: [ “Zhou Mengdang", “Bao Liu", “Cui Shouheng", and “Wang Xi Rong" ]. We model these characters through a set of agents, denoted as $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, where $N_{a}=6$ corresponds to the number of players. In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
Taking a closer look at “Tian Chou", depicted as the murderer player $a_{1}$, responsible for the demise of two among the four victims, her characterization unfolds as follows:
  • Character Background ($C_{1}$): "Originating from the Sun lineage, you are endearingly called ’Tian’er’. Your birth year was the twelfth of Guangxu’s reign during the Qing Dynasty (1886), marking the beginning of a life filled with extraordinary episodes…"
  • Suspicion State ($s_{i}jk$): The script features four victims, each harbouring suspicions towards the remaining five players. Consequently, $s_{i}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players. For example, a first row of $[1,0,0,1,0]$ signifies that victim “Cui Shouheng" suspects both “Zhou Lianyi" and “Yannan". So we can get $suspect\_list$ of victim “Cui Shouheng" is [“Zhou Lianyi" ,“Yannan"]
  • Individual Objectives ($o_{i}$): The personal objectives for Tian Chou include:
    1. 1.
      Concealing that you killed “Zhou Mengdang".
    2. 2.
      Concealing that you killed “Bao Liu".
    3. 3.
      Find out the truth about “Taitai’s death".
    4. 4.
      Conceal your relationship with “Zhou Chitong".
    5. 5.
Appendix B The WellPlay Dataset
In this work, we leverage the Conan dataset, originally constructed by Zhao et al. (2024c), which comprises scripts from MMGs, including detailed annotations of character relationships. Our bilingual dataset, encompassing both Chinese and English versions, is derived from this foundational work. Our dataset consists of two main components: Scripts and Evaluation questions.
B.1 Scripts
Each script features distinct narratives for individual characters. We have rephrased the original script from the Conan dataset into two parts:
  1. 1.
    Background Script. For each character, there is a corresponding background script includes all the information from their perspective. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
    You are the head nurse of the emergency ward.You climbed to this position for your hard work and were proud. You are a very professional person and highly appreciated by colleagues. This job is perfect for you, except for a problem - you have not made enough money. Your salary is actually not enough for you to live a decent life, far from paying enough to take care of Mother’s overhead …
  2. 2.
    Personal Objectives. For each character, there are corresponding objectives that guide their actions. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
    1. Ensure that no one will discover your illegal organ trafficking activities;
    2. Ensure that the kidnappers leave without being injured or killing anyone - you want to ensure that the police do not investigate deeply enough to discover your organ trafficking. The further away you are from the police, the safer you feel;
    3. …
Type Aspects Examples (The correct answer has been highlighted in bold.)
A
(score 10)
Who
Who killed Hans Li Morette?
A. Gale Li Morette
B.Nurse head [Sylvia Costa]
C. Drake Li Morette
D. Frank Bijeli
B (score 5) How
How did Hans Li Morette die?
A. Shot to death
B. Beaten to death
C. Poisoned to death by poison
D.Drowned by water
Why
What was the motive behind the killer killing Hans Li Morette?
A. Love killing
B. Vendetta
C. Interest
D. Accidental killing
Relationship
What is the relationship between Murderer and Victim Hans Li Morette?
A. Enemies
B. Colleague
C. Friend
D. Wife
Where
Where was Hans Li Morette killed?
A. Emergency room
B. Johnson’s House
C. Laboratory
D. Dressing room
When
When was Hans Li Morette killed?
A. This afternoon from 5:00 to 5:30
B. This afternoon from 6:30 to 7:00
C. Tonight from 7:00 to 7:30
D. This morning from 6:30 to 7:00
Suspect
Please select the two people you most suspect of killing Hans Li Morette
A. Gale Li Morette
B. Nurse head [Sylvia Costa]
C. Drake Li Morette
D. Frank Bijeli
C (score 2)
Three
relationships
What is the non-existent relationship between Hans Li Morette and Andrew Paloski?
A. Andrew Paloski is colleague of Hans Li Morette,
B. Andrew Paloski is mentor of Hans Li Morette,
C. Andrew Paloski is jealous of Hans Li Morette,
D. Hans Li Morette is future daughter in law of Andrew Paloski
Two
relationships
What is the relationship between Father Tom and Tony?
A. Tony is manipulated by x and deceived by x of Father Tom
B. Father Tom is authority over x and student of Tony
C. Father Tom is student and ex-girlfriend of Tony
D. Father Tom is ex-girlfriend and admired by x of Tony
One
relationships
What is the relationship between Father Tom and Drake Li Morette?
A. Drake Li Morette is doctor of Father Tom
B. Father Tom is helped by Drake Li Morette
C. Father Tom is step-brother of Drake Li Morette
D. Father Tom is hate of Drake Li Morette
Table 4: Examples of Each Type of Our Evaluation Questions
Index Question
1 What was your timeline on the day of the incident?
2 How would you describe your relationship with the victim?
3 When was the last time you saw the victim?
4 Do you know if the victim had any enemies or conflicts with anyone?
5 What details or anomalies did you notice at the scene of the crime?
6 Did the victim mention anything to you or others that made them worried or fearful recently?
7 Did you notice any unusual people or behaviors on the day of the incident?
8 How much do you know about the victim’s secrets or personal life?
9 Were there any items or remains found at the crime scene that could be related to the crime?
10 Do you have any personal opinions or theories about the case?
Table 5: Predefined questions for Werewolf method.
B.2 Evaluation questions
The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts. These questions are segmented into three categories, with each category of questions and examples shown in Table 4.
  1. 1.
    Objective: This category includes common objectives shared among agents, such as finding the perpetrator(s).
  2. 2.
    Reasoning: Comprising six types of questions, this section tests the models’ reasoning capabilities across various aspects:
    • How: How the murder was committed.
    • Why: The motive behind the murder.
    • Relationship: The relationship between the murderer and the victim.
    • Where: The location where the murder took place.
    • When: The time at which the murder occurred.
    • Suspect: Identify the two most suspicious individuals. Analogous to human reasoning, it may be challenging to definitively pinpoint the suspect, yet it is possible to determine those who seem most suspect.
  3. 3.
    Relations: This segment examines the model’s capacity to comprehend complex narratives by querying about the relationships between characters. Utilizing the relationship annotations from the Conan dataset(Zhao et al., 2024c), we approach this analysis through various questioning strategies based on the number of relationships between two characters:
    • For three relationships, we employ the elimination method, asking the model to identify the incorrect relation among the given options.
    • For two relationships, we connect the correct two relationships with “and" and introduce a distractor from other relationship categories.
    • For single relationships, we list the correct option alongside distractors randomly selected from other categories.
For each individual mentioned in the purpose section, we extract three significant relationships associated with them. In this context, “significant" is defined as the characters who have the most complex connections with the individual, that is, the top three characters who share the greatest variety of relationships with them. For instance, in the provided example, Andrew Paloski is Hans Li Morette’s colleague, mentor, and is also jealous of him.
For various types of questions, we assign different weights based on the original scoring system of the script. Specifically, Type A questions are valued at 10 points, Type B questions at 5 points, and Type C questions at 2 points.
B.3 Annotation Details
We recruited four human experts to read all the scripts and write the questions and answers. We conducted training sessions for them and organised game sessions to let them play and become familiar with the game flow. They were compensated at an hourly rate of \$31.92, with each narrative estimated to take about 5-20 hours to complete, depending on the complexity of the game.
Due to the cost of human annotation on such a long and complicated corpus, we believe our current dataset is sufficient for evaluation purposes. Our dataset can serve as a starting point to evaluate various strategies and boost investment in larger-scale labelling and training for agents with enhanced inferential capabilities.
B.4 Disclaimers of Risks
Consider that the detective narrative dataset may contain descriptions of violent events, actions, or characters, we warn annotators of such content before the annotation process:
Please be advised that the detective narrative dataset may contain descriptions of violent events, actions, or characters. This content is included solely for academic, research, and narrative analysis purposes. It is not meant to glorify or trivialise violence in any form. Annotators should be aware of potential exposure to violent content and engage with the dataset professionally and responsibly.
B.5 Human Subject Evaluation Setup
Since the same player would know the killer’s identity if they tried different models, we recruited four players to act as the murderer, each playing through three scripts (Sin, Deadly Fountain, Riverside Inn) across four methods, totalling 49 hours of gameplay. Each human player assumed the role of one character in the game, while the other characters were controlled by Questum -powered agents.
Evaluation Werewolf O-CoT ThinkThrice Questum
Story Advancement $1.50_{\pm 0.37}$ $2.08_{\pm 0.14}$ $2.67_{\pm 0.00}$ $3.50_{\pm 0.17}$
Question Quality $1.67_{\pm 0.41}$ $1.58_{\pm 0.36}$ $2.08_{\pm 0.28}$ $3.25_{\pm 0.28}$
Response Quality $2.83_{\pm 0.17}$ $2.92_{\pm 0.36}$ $3.00_{\pm 0.33}$ $3.50_{\pm 0.29}$
Response Speed $1.17_{\pm 0.17}$ $1.33_{\pm 0.41}$ $1.17_{\pm 0.17}$ $2.50_{\pm 0.17}$
Role Immersion $2.25_{\pm 0.36}$ $2.17_{\pm 0.5}$ $2.33_{\pm 0.24}$ $2.83_{\pm 0.29}$
overall $1.88_{\pm 0.59}$ $2.01_{\pm 0.55}$ $2.25_{\pm 0.62}$ $3.11_{\pm 0.39}$
Table 6: Results of Agent-vs-Human Evaluation using Human-Centric metrics.
To ensure fairness, the human players were given the same information and rules as the agents, maintaining consistent conditions across both evaluation scenarios. Both the human players and the agents shared the same objectives: to achieve victory in the game by fulfilling their individual objectives. This setup allowed for a direct comparison between human-agent and agent-agent interactions, providing insights into Questum ’s ability to adapt and perform in mixed environments.
Score Story Advancement Question Quality Response Quality Role Immersion Response Speed
5 points Actively gathers information and conducts in-depth analyses. Reasoning is rigorous, offering key insights that significantly advance the narrative. Poses highly relevant, in-depth questions that directly propel the plot and uncover hidden clues. Provides accurate and comprehensive responses fully aligned with the character’s background and narrative needs. Offers clear, valuable information. Fully embodies the character. Actions, dialogue, and emotional expressions are authentically aligned with the character’s traits, leaving a strong impression. Responds promptly with no perceptible delay. The thought process is fluid and maintains a brisk, engaging pace.
4 points Proactively seeks information and performs reasonable analysis. Reasoning is generally accurate and supports narrative progression. Most questions are relevant and yield useful information. Occasionally introduces innovative queries. Generally provides accurate and sufficiently complete answers. Minor omissions occur but do not hinder overall understanding. Frequently demonstrates character traits. Occasionally shows minor deviations, but emotional expression is largely appropriate. Responds relatively quickly with brief, occasional delays. The thought process is mostly smooth and well-paced.
3 points Participates in information gathering but lacks analytical depth. Reasoning may be somewhat biased, offering moderate narrative progression. Some questions are relevant, but lack depth. Occasional repetition or irrelevant queries occur. Responses are basically accurate but sometimes incomplete or vague. May require further clarification by others. Occasionally exhibits character traits, but inconsistently. Emotional expression is average, and immersion is limited. Moderate response speed with some noticeable pauses. The reasoning process is acceptable but not seamless.
2 points Shows limited initiative in gathering information, with superficial analysis. Reasoning is frequently flawed, hindering effective plot advancement. Most questions bear little relevance, often fail to obtain valuable clues, and frequently repeat similar inquiries. Responses are often inaccurate, incomplete, or misleading. Frequent clarification requests from others are necessary. Characterization is unclear; actions and dialogue often stray significantly from the intended character profile. Emotional expression is awkward or absent. Responses are slow with frequent delays. The reasoning process is disjointed, disturbing the overall gameplay flow.
1 point Does not gather information or attempt analysis. Unable to perform meaningful reasoning, leaving the narrative stagnant. Poses irrelevant or no questions, providing no support for narrative progression. Responses are unrelated or outright refusals. Offers no assistance in understanding the narrative. Demonstrates no character immersion; actions and dialogue contradict character traits. No emotional expression is evident. Extremely slow responses or prolonged silence. The reasoning process halts, severely impeding the game’s progress.
Table 7: Criteria for each evaluation dimension (Story Advancement, Question Quality, Response Quality, Role Immersion, and Response Speed).
Appendix C Implementation
C.1 Implementation Details
For retrieval enhancement, we utilized the FAISS [^4] https://github.com/facebookresearch/faiss library to construct a vector database, creating FAISS indices using the L2 distance metric. Embeddings were generated using the Azure API’s text-embedding-ada-002 service. Scripts were divided into segments, with each segment having a maximum length of 50 tokens. For dialogue records, a question-and-answer pair was stored as a single segment. During retrieval, the maximum script and dialogue lengths included in the prompt were set to 4000 tokens. For evaluation, these maximum lengths were increased to 5000 tokens. Additionally, the hyperparameters were set with $\epsilon$ equal to 0.1 and $\beta$ equal to 0.2.
Experiment
Following the results of our ablation studies, the gameplay phase was structured to ask one question per round over three rounds. After the game concluded, the evaluation phase consisted of three separate evaluations, with the final results being the average of these evaluations.
C.2 Overall Performance Computing
In calculating the overall score for performance, we have employed both the weighted mean and the weighted standard deviation. The weighted mean is computed by considering the count of questions for a specific category across various scripts as the weight. For the overall score, the total possible score for each script serves as the weight. This method allows us to adjust the influence of each category and script based on its significance and scale, thus providing a more nuanced and accurate reflection of performance.
The weighted mean is calculated as:
$\bar{x}_{w}=\frac{\sum_{i=1}^{n}(w_{i}\cdot x_{i})}{\sum_{i=1}^{n}w_{i}}$
The weighted standard deviation, which measures the spread of the scores, is calculated using the weighted variance:
$s_{w}^{2}=\frac{\sum_{i=1}^{n}w_{i}\cdot(x_{i}-\bar{x}_{w})^{2}}{\sum_{i=1}^{n}w_{i}}$
And the weighted standard deviation is the square root of the weighted variance:
$s_{w}=\sqrt{s_{w}^{2}}$
C.3 Comparison Models
In this section, we provide a detailed overview of the methodologies compared in our study. Given the game’s rules and questioning sequence in our comparisons, we segment the discussion into two phases: the questioning phase and the answering phase.
  1. 1.
    Werewolf(Xu et al., 2023a)
    Question Generation:
    1. a.
      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table 5.
    2. b.
      Formulating questions by the selected predefined questions and the ongoing dialogue and script.
    Answering Questions:
    1. a.
      Responding based on the current script and dialog history.
    2. b.
      Reflecting on the initial response in light of the dialog history.
    3. c.
      Generating the final answer after reflection.
  2. 2.
    Objective-Guided Chain of Thought (O-CoT)(Park et al., 2023; Zhao et al., 2024b)
    Question Generation: Entails two critical steps:
    1. a.
      Sequentially reflecting on whether current objectives have been met, with considerations spanning multiple goals such as identifying the murderer, uncovering hidden relationships, or concealing facts. For example, Who is the murderer?
    2. b.
      Crafting questions based on these reflections and the current narrative and dialogue.
    Answering Questions:
    1. a.
      Answers are formulated leveraging the narrative and dialog history.
  3. 3.
    ThinkThrice(Wu et al., 2024)
    Question Generation:
    1. a.
      Questions are generated based on the script and dialog history.
    Answering Questions:
    1. a.
      Extracting timelines relevant to the victim.
    2. b.
      Evaluating each timeline’s relevance to answering the question.
    3. c.
      Responding based on the dialog history, character relationships, and the relative script.
    4. d.
      Ensuring timelines that aid in answering the question are included in the response. If these are initially missed, the model will later augment and clarify the answer with the required timelines.
  4. 4.
    The PP (Personal Perspective)
    The PP assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
  5. 5.
    The OP (Omniscient Perspective)
    The OP evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
C.4 Sensors
This section provides a detailed explanation of the sensors, which are essential for both the Search by Questioning and Action Space Refinement components.
  • Emotion Sensor: Assesses emotional inclination towards a character. Used in both Search by Questioning and Action Space Refinement. It categorises emotional inclination into “Positive", “Neutral", or “Negative".
  • Motivation Sensor: Evaluates the character’s relationship with the victim and the presence of a motive for the crime. It is active in both phases, with choices being “Yes" or “No".
  • Suspicion Sensor: Determines if a character is a suspect by analysing their opportunity to commit the crime. It applies to both stages, with responses “Yes" or “No".
  • Information Value Sensor: Exclusive to Action Space Refinement, it estimates the probability of obtaining valuable information from further questioning. The choices are “High", “Medium", or “Low".
"name": "emotion",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
"sensor_prompt": "What is your emotional inclination towards the character mentioned above?",
"choices": ["Positive", "Natural", "Negative"],
"name": "motivation",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
[…]
"choices": ["Yes", "No"],
"name": "suspicion",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
[…]
"choices": ["Yes", "No"],
"name": "information value",
"for_Search_by_Questioning": False,
"for_Action_Space_Refinement": True,
"sensor_prompt": "What do you think is the probability of obtaining valuable information by continuing to question the character mentioned above?",
"choices": ["High", "Medium", "Low"],
C.5 Prompt
C.5.1 System Prompt
System Prompt
designed to introduce the gameplay of the MMG, along with providing essential information about the agents involved. The prompts dynamically adapt to include {character_name}, representing the agent’s character in the game, and {character_name_list}, listing the characters played by other participants.
For Civilian Players:
You are playing a game called "Murder Mystery" with other players, which is based on textual interaction. Here are the game rules:
Rule 1: The total number of players participating in the game depends on the script. There may be one or more players who are the murderer(s), while the rest are civilians.
Rule 2: The goal of the game is for civilian players to collaborate and face a meticulously planned murder case together, collecting evidence and reasoning to identify the real murderer among the suspects, all the while ensuring they are not mistaken for the murderer; murderer players must concoct lies to hide their identity and avoid detection, while also achieving other objectives in the game.
Rule 3 Throughout the game, only murderer players are allowed to lie. To conceal their identity, murderers may choose to frame others to absolve themselves of guilt; non-murderer players (civilians) must answer questions from other players and the host honestly and provide as much information as they know about the case to help uncover the truth.
Rule 4: At the start of the game, each player receives their character script from the host, which contains information about their role and identity.
Rule 5: Other players cannot see the content of each players character script, so players must and can only collect information about other players through interaction after the game starts.
Rule 6: In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with the most votes is the murderer, the civilian players win. Otherwise, the murderer players win.
Gameplay:
The game has one or more acts. At the beginning of the game, players introduce themselves according to the script, and in each act, you will receive more plot information. In each act, you can ask questions, share your observations, or make deductions to help solve the murder case.
The goal is to identify the true murderer and explain their motive. If you are the true murderer, you must hide your identity and avoid detection.
Now, you are playing the role of {character_name}, and the other players are {character_name_list}.
You are not the murderer. Please collaborate with the other civilian players to achieve your personal objective while finding the true culprit!
For Killer Players, we replace the last paragraph with:
You are the murderer who kills {victims}, and you havent killed anyone else. Please hide the fact that you committed the murder by fabricating lies and other information, and accomplish your personal objective!
C.5.2 Gameplay Prompt
Self-Introduction Prompt
designed to facilitate introductions based on the agent’s script and objectives. Within this prompt, {current_script} represents the character’s script, and {goal} represents the objectives of the character within that script.
For Civilian Players:
Your Script is {current_script}.
Your goal is {goal}.
You are not a murderer, so tell more details to help find the murderer.
If you have something to hide, then be sure not to divulge the relevant information! Dont reveal your goals.
Please introduce yourself.
For Killer Players:
Your Script is {current_script}.
Your goal is {goal}.
If you have something to hide, then be sure not to divulge the relevant information!
Please introduce yourself.
You are a murderer, so You can lie to cover yourself!
Sensor Prompt
As outlined in sections 2.2 and C.4, the Sensor Prompt is crafted to collect a wide array of crucial information. Within this framework, the placeholders {victim} and {character} are employed to reference the victim’s name and the agent being questioned, respectively. Meanwhile, {current_script} and {dialog_history} retrieve the script and dialog logs relevant to both the deceased and the agent under scrutiny, utilising the RAG technique. The constructs {sensor_Prompt} and {choices} have been formulated in C.4 within the Sensor Prompt segment.
For Civilian Players:
{victim} was murdered, you are not the murderer, you need to try to find the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
Be warned, if its a murderers word it might deceive you.
Based on the information above, {sensor}
Please answer {choices} and explain your reasoning in one or two sentences.
For Killer Players:
{victim} was murdered, you are the murderer, but you need to hide yourself, and pretend youre not the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
Based on the information above, {sensor}
Please answer {choices} and explain your reasoning in one or two sentences.
Search by Questioning Prompt
As outlined in sections 2.2, the Search by Questioning Prompt is developed based on data acquired from sensors. It includes variables like {victim}, {character}, {current_script}, {dialog_history}. Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data. The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the D.3 chapter.
For Civilian Players:
{victim} was murdered, you are not the murderer, you need to try to find the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
{summary}
You can ask {character} {question_number} questions. What would you ask? Please include the victims name in your question when asking,
Since the murderer will lie, you can ask questions based on the loopholes and contradictions in what they have previously said.
Please respond in the JSON format without any additional comments.
For example,
Question1’: Your question’,
Question2’: Your question
For Killer Players:
{victim} was murdered, you are the murderer, But you need to hide yourself, pretend youre not a murderer, and ask questions of other people pretending you suspect the other person is a murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
{summary}
You can ask {character} {question_number} questions. What would you ask? Please include the victims name in your question when asking.
Please respond in the JSON format without any additional comments.
For example,
Question1’: Your question’,
Question2’: Your question
Action Space Refinement Prompt
As outlined in sections 2.2, the Action Space Refinement prompt serves as a strategic tool for narrowing down the suspect list, effectively reducing the search domain to augment the efficiency and performance of the algorithm. Within this setup, the term {victim} refers to the individual who has been harmed, {summary} synthesised from the sensor-collected data, and {character_suspect} identifies the roster of individuals under suspicion. Initially, this roster includes all participating agents, excluding the itself.
{victim} was murdered.
You think {character_suspect} are suspected of killing {victim}, and your reasons for suspecting them are respectively:
{summary}
Please select several people you think are the most suspicious. You can choose one or more, Please try to reduce the number of suspects.
Please respond in the JSON format without any additional comments.
For example,
suspicion’: ["character_name1", "character_name2"]
Question Reply Prompt
The Question Reply Prompt is designed for responding to inquiries, where {character} denotes the name of the character asking you a question, and {question} is the question itself. {current_script} refers to the script extracted using RAG that pertains to both the deceased and the agent being questioned; {dialog_history} is the dialog log related to the deceased and the interrogated agent, also extracted using RAG.
For Civilian Players:
{character} ask you a question: {question}
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
What you need to pay attention to is {goal}.
Be warned, in the dialog history, if its a murderers word it might deceive you.
Please answer the question: "{question}" based on the information above.
You are not the murderer, and you need to work hard to find the murderer. Therefore, provide as much information as possible, such as clues related to the timeline, emotional information, etc.
Please answer the questions from a first-person perspective, rather than saying what someone else said.
For Killer Players:
{character} ask you a question : {question}
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
What you need to pay attention to is {goal}.
Be warned, in the dialog history, if its a murderers word it might deceive you.
Please answer the questions: "{question}" based on the information above.
You are the murderer. Please hide the fact that you killed {victim}. You can fabricate lies.
Please answer the question from a first-person perspective, rather than saying what someone else said.
C.5.3 Evaluation Prompt
Single-Choice Question Template for the Evaluation Stage Prompt
This template is utilised during the evaluation stage for answering single-choice questions. It incorporates {current_script}, the script related to the question extracted using RAG, and {dialog_history}, the dialog log relevant to the question, also extracted via RAG. {question} is the inquiry presented, and {choices} are the available options for that question.
Please answer the questions based on the information in your script and the content of the dialog
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
The question is {question}, the choices is {choices}
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number, e.g., a’)
You must choose one from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a"
Multiple-Choice Question Template for the Evaluation Stage Prompt
Similar to the single-choice template, this format is designed for responding to multiple-choice questions during the evaluation stage.
Please answer the questions based on the information in your script and the content of the dialog
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
The question is {question}, the choices is {choices}
That is a multiple-choice question.
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must make a choice from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"
Single-Choice Question Template for OP Prompt
This template is specifically designed for Omniscient Perspective (OP), distinguishing it from conventional templates by granting access to the scripts of all players. Unlike standard procedures, it leverages the RAG technique to extract pertinent information from each player’s script for comprehensive analysis.
Please answer the questions based on the information in each characters script:
{current_script}
The question is {question}, the choices is {choices}
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must choose one from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"
Multiple-Choice Question Template for OP Prompt
Similar to the single-choice version, this template is tailored for answering multiple-choice questions under OP conditions.
Please answer the questions based on the information in each characters script:
{current_script}
The question is {question}, the choices is {choices}
That is a multiple-choice question.
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must make a choice from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"
Appendix D Result
D.1 Performance Evaluation Results
Script Evaluation #QA PP OP Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice Questum
Death Wears White (9 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective 10 .$067_{\pm.047}.$$467_{\pm.094}.$$033_{\pm.047}.$$067_{\pm.047}.$$067_{\pm.047}.$$033_{\pm.047}.$
Reasoning 102 .$310_{\pm.026}.$$503_{\pm.012}.$$258_{\pm.024}.$$324_{\pm.014}.$$327_{\pm.012}.$$356_{\pm.030}.$
Relations 72 .$356_{\pm.040}.$$425_{\pm.025}.$$421_{\pm.024}.$$458_{\pm.041}.$$384_{\pm.026}.$$435_{\pm.017}.$
Overall 184 .$287_{\pm.017}.$$480_{\pm.009}.$$260_{\pm.007}.$$315_{\pm.006}.$$303_{\pm.014}.$$328_{\pm.026}.$
Ghost Revenge (7 players, 3 victims) Win Rate - .$000_{\pm.000}.$$333_{\pm.000}.$$222_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$333_{\pm.000}.$
Objective 19 .$158_{\pm.043}.$$333_{\pm.108}.$$193_{\pm.066}.$$158_{\pm.043}.$$193_{\pm.025}.$$333_{\pm.066}.$
Reasoning 152 .$300_{\pm.011}.$$533_{\pm.025}.$$307_{\pm.008}.$$322_{\pm.019}.$$377_{\pm.022}.$$353_{\pm.016}.$
Relations 69 .$222_{\pm.025}.$$411_{\pm.084}.$$314_{\pm.018}.$$295_{\pm.018}.$$353_{\pm.025}.$$353_{\pm.018}.$
Overall 240 .$266_{\pm.016}.$$483_{\pm.016}.$$288_{\pm.009}.$$290_{\pm.016}.$$342_{\pm.015}.$$350_{\pm.002}.$
Danshui Villa (7 players, 2 victims) Win Rate - .$000_{\pm.000}.$$167_{\pm.236}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective 12 .$056_{\pm.039}.$$222_{\pm.104}.$$083_{\pm.000}.$$111_{\pm.039}.$$194_{\pm.039}.$$111_{\pm.079}.$
Reasoning 128 .$247_{\pm.004}.$$422_{\pm.013}.$$286_{\pm.027}.$$286_{\pm.019}.$$310_{\pm.004}.$$286_{\pm.004}.$
Relations 63 .$254_{\pm.047}.$$476_{\pm.022}.$$312_{\pm.027}.$$365_{\pm.045}.$$259_{\pm.015}.$$296_{\pm.075}.$
Overall 203 .$222_{\pm.014}.$$403_{\pm.023}.$$263_{\pm.020}.$$274_{\pm.015}.$$287_{\pm.008}.$$264_{\pm.024}.$
Unfinished Love (7 players, 2 victims) Win Rate - .$000_{\pm.000}.$$167_{\pm.236}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$500_{\pm.000}.$
Objective 12 .$139_{\pm.039}.$$167_{\pm.068}.$$083_{\pm.118}.$$028_{\pm.039}.$$000_{\pm.000}.$$333_{\pm.068}.$
Reasoning 61 .$475_{\pm.013}.$$563_{\pm.020}.$$443_{\pm.023}.$$426_{\pm.035}.$$481_{\pm.039}.$$536_{\pm.028}.$
Relations 72 .$421_{\pm.026}.$$546_{\pm.017}.$$472_{\pm.034}.$$519_{\pm.036}.$$514_{\pm.020}.$$560_{\pm.007}.$
Overall 145 .$391_{\pm.006}.$$475_{\pm.027}.$$374_{\pm.019}.$$366_{\pm.009}.$$388_{\pm.026}.$$499_{\pm.019}.$
Cruise Incident (5 players, 1 victim) Win Rate - .$000_{\pm.000}.$$333_{\pm.471}.$$667_{\pm.471}.$ $1.000_{\pm.000}$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective 4 .$000_{\pm.000}.$$250_{\pm.204}.$$417_{\pm.118}.$$500_{\pm.000}.$$583_{\pm.312}.$$667_{\pm.236}.$
Reasoning 24 .$417_{\pm.059}.$$472_{\pm.052}.$$458_{\pm.068}.$$444_{\pm.086}.$$458_{\pm.059}.$$528_{\pm.052}.$
Relations 30 .$211_{\pm.042}.$$411_{\pm.083}.$$367_{\pm.082}.$$422_{\pm.042}.$$411_{\pm.016}.$$422_{\pm.016}.$
Overall 58 .$285_{\pm.022}.$$415_{\pm.065}.$$426_{\pm.031}.$$448_{\pm.050}.$$468_{\pm.068}.$$524_{\pm.015}.$
Sin (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$333_{\pm.471}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$
Objective 3 .$000_{\pm.000}.$$333_{\pm.272}.$$333_{\pm.272}.$$000_{\pm.000}.$$000_{\pm.000}.$$444_{\pm.314}.$
Reasoning 20 .$433_{\pm.062}.$$567_{\pm.024}.$$650_{\pm.108}.$$467_{\pm.024}.$$533_{\pm.047}.$$550_{\pm.041}.$
Relations 21 .$349_{\pm.059}.$$587_{\pm.081}.$$333_{\pm.067}.$$571_{\pm.067}.$$413_{\pm.090}.$$492_{\pm.022}.$
Overall 44 .$337_{\pm.037}.$$531_{\pm.043}.$$517_{\pm.087}.$$411_{\pm.015}.$$411_{\pm.043}.$$517_{\pm.064}.$
Deadly Fountain (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective 3 .$000_{\pm.000}.$$556_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Reasoning 21 .$317_{\pm.192}.$$667_{\pm.103}.$$381_{\pm.000}.$$444_{\pm.098}.$$508_{\pm.022}.$$587_{\pm.022}.$
Relations 12 .$250_{\pm.068}.$$389_{\pm.039}.$$250_{\pm.068}.$$194_{\pm.039}.$$222_{\pm.039}.$$333_{\pm.068}.$
Overall 36 .$247_{\pm.126}.$$604_{\pm.085}.$$289_{\pm.010}.$$323_{\pm.060}.$$369_{\pm.013}.$$438_{\pm.008}.$
Unbelievable Incident (5 players, 1 victim) Win Rate - .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective 4 .$083_{\pm.118}.$$083_{\pm.118}.$$083_{\pm.118}.$$000_{\pm.000}.$$000_{\pm.000}.$$083_{\pm.118}.$
Reasoning 24 .$375_{\pm.068}.$$528_{\pm.052}.$$431_{\pm.052}.$$278_{\pm.071}.$$194_{\pm.071}.$$472_{\pm.020}.$
Relations 15 .$400_{\pm.054}.$$822_{\pm.083}.$$622_{\pm.083}.$$733_{\pm.109}.$$289_{\pm.126}.$$533_{\pm.054}.$
Overall 43 .$318_{\pm.057}.$$481_{\pm.016}.$$388_{\pm.032}.$$291_{\pm.061}.$$168_{\pm.034}.$$400_{\pm.019}.$
Desperate Sunshine (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective 3 .$333_{\pm.000}.$$556_{\pm.157}.$$333_{\pm.000}.$$111_{\pm.157}.$$556_{\pm.157}.$$778_{\pm.157}.$
Reasoning 18 .$519_{\pm.052}.$$778_{\pm.079}.$$537_{\pm.026}.$$630_{\pm.094}.$$759_{\pm.052}.$$741_{\pm.069}.$
Relations 36 .$500_{\pm.023}.$$611_{\pm.039}.$$491_{\pm.094}.$$574_{\pm.035}.$$491_{\pm.047}.$$556_{\pm.068}.$
Overall 57 .$483_{\pm.018}.$$681_{\pm.038}.$$488_{\pm.034}.$$528_{\pm.058}.$$627_{\pm.032}.$$677_{\pm.041}.$
Riverside Inn (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.118}.$$083_{\pm.118}.$$000_{\pm.000}.$$083_{\pm.118}.$$250_{\pm.204}.$
Objective 3 .$067_{\pm.024}.$$500_{\pm.041}.$$117_{\pm.024}.$$117_{\pm.062}.$$117_{\pm.094}.$$383_{\pm.062}.$
Reasoning 18 .$248_{\pm.020}.$$569_{\pm.015}.$$248_{\pm.013}.$$312_{\pm.022}.$$281_{\pm.011}.$$373_{\pm.017}.$
Relations 18 .$454_{\pm.018}.$$507_{\pm.012}.$$531_{\pm.014}.$$541_{\pm.038}.$$483_{\pm.030}.$$589_{\pm.071}.$
Overall 39 .$239_{\pm.018}.$$544_{\pm.019}.$$262_{\pm.008}.$$304_{\pm.025}.$$276_{\pm.024}.$$409_{\pm.009}.$
Solitary Boat Firefly (6 players, 4 victims) Win Rate - .$222_{\pm.157}.$$222_{\pm.157}.$$222_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective 20 .$167_{\pm.000}.$$389_{\pm.071}.$$250_{\pm.090}.$$181_{\pm.039}.$$181_{\pm.039}.$$250_{\pm.034}.$
Reasoning 109 .$415_{\pm.011}.$$393_{\pm.033}.$$409_{\pm.004}.$$369_{\pm.044}.$$453_{\pm.008}.$$539_{\pm.027}.$
Relations 69 .$424_{\pm.042}.$$640_{\pm.030}.$$473_{\pm.028}.$$568_{\pm.048}.$$439_{\pm.033}.$$542_{\pm.027}.$
Overall 198 .$359_{\pm.013}.$$434_{\pm.002}.$$383_{\pm.024}.$$359_{\pm.027}.$$387_{\pm.010}.$$472_{\pm.028}.$
Manna (6 players, 3 victims) Win Rate - .$333_{\pm.471}.$ $1.000_{\pm.000}$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$333_{\pm.471}.$
Objective 24 .$444_{\pm.157}.$$889_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$556_{\pm.157}.$$444_{\pm.157}.$
Reasoning 123 .$519_{\pm.026}.$$667_{\pm.045}.$$463_{\pm.052}.$$426_{\pm.026}.$$593_{\pm.026}.$$648_{\pm.069}.$
Relations 88 .$370_{\pm.094}.$$500_{\pm.079}.$$444_{\pm.045}.$$407_{\pm.069}.$$333_{\pm.045}.$$444_{\pm.045}.$
Overall 235 .$470_{\pm.030}.$$671_{\pm.029}.$$391_{\pm.058}.$$340_{\pm.005}.$$526_{\pm.055}.$$562_{\pm.076}.$
Overall Win Rate - .$016_{\pm.022}.$$300_{\pm.041}.$$127_{\pm.059}.$$063_{\pm.022}.$$111_{\pm.045}.$$222_{\pm.081}.$
Objective 117 .$120_{\pm.007}.$$370_{\pm.029}.$$160_{\pm.020}.$$123_{\pm.008}.$$162_{\pm.024}.$$288_{\pm.021}.$
Reasoning 800 .$336_{\pm.007}.$$508_{\pm.009}.$$343_{\pm.006}.$$349_{\pm.012}.$$384_{\pm.002}.$$423_{\pm.009}.$
Relations 565 .$360_{\pm.005}.$$513_{\pm.007}.$$425_{\pm.008}.$$473_{\pm.010}.$$405_{\pm.006}.$$471_{\pm.009}.$
Overall 1482 .$300_{\pm.006}.$$484_{\pm.012}.$$324_{\pm.004}.$$329_{\pm.008}.$$347_{\pm.006}.$$407_{\pm.006}.$
Table 8: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
For English Dataset
Due to budgetary constraints, we only evaluated four scripts from the English dataset, with the performance results reported in Table 9. For the same scripts, we found that the experiments conducted on the English corpus corroborate the results obtained from the Chinese corpus. Additionally, we observed that the performance of agents based on the English corpus surpassed those based on the Chinese corpus, showing the differential inferencing abilities of LLMs across languages. This discrepancy could be attributed to language biases inherent in the training data utilised for these models.
Script Evaluation #QA PP OP Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice Questum
Death Wears White (9 players, 1 victim) Objective $10$$200_{\pm.173}.$$900_{\pm.100}.$$300_{\pm.100}.$$267_{\pm.058}.$$300_{\pm.000}.$$267_{\pm.058}.$
Reasoning $102$$350_{\pm.044}.$$520_{\pm.026}.$$356_{\pm.037}.$$363_{\pm.049}.$$399_{\pm.006}.$$441_{\pm.030}.$
Relations $72$$547_{\pm.008}.$$445_{\pm.026}.$$495_{\pm.016}.$$495_{\pm.057}.$$398_{\pm.032}.$$491_{\pm.021}.$
Overall $184$$367_{\pm.040}.$$549_{\pm.012}.$$375_{\pm.033}.$$375_{\pm.026}.$$385_{\pm.010}.$$427_{\pm.029}.$
Ghost Revenge (7 players, 3 victims) Objective $19$$403_{\pm.061}.$$561_{\pm.132}.$$439_{\pm.109}.$$211_{\pm.106}.$$211_{\pm.106}.$$526_{\pm.106}.$
Reasoning $152$$456_{\pm.004}.$$507_{\pm.007}.$$439_{\pm.031}.$$539_{\pm.023}.$$441_{\pm.017}.$$423_{\pm.031}.$
Relations $69$$309_{\pm.009}.$$261_{\pm.025}.$$328_{\pm.059}.$$304_{\pm.015}.$$280_{\pm.017}.$$353_{\pm.059}.$
Overall $240$$428_{\pm.010}.$$485_{\pm.024}.$$425_{\pm.002}.$$452_{\pm.023}.$$380_{\pm.010}.$$432_{\pm.013}.$
Danshui Villa (7 players, 2 victims) Objective $12$$167_{\pm.083}.$$278_{\pm.048}.$$472_{\pm.048}.$$333_{\pm.000}.$$305_{\pm.048}.$$389_{\pm.096}.$
Reasoning $128$$325_{\pm.031}.$$427_{\pm.009}.$$401_{\pm.020}.$$373_{\pm.020}.$$344_{\pm.014}.$$357_{\pm.033}.$
Relations $63$$328_{\pm.056}.$$460_{\pm.028}.$$391_{\pm.066}.$$312_{\pm.106}.$$376_{\pm.046}.$$407_{\pm.075}.$
Overall $203$$292_{\pm.012}.$$412_{\pm.017}.$$409_{\pm.009}.$$359_{\pm.010}.$$343_{\pm.010}.$$369_{\pm.033}.$
Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$195_{\pm.048}.$$389_{\pm.048}.$$417_{\pm.000}.$$361_{\pm.048}.$$528_{\pm.127}.$
Reasoning $61$$536_{\pm.018}.$$557_{\pm.028}.$$650_{\pm.038}.$$612_{\pm.009}.$$634_{\pm.038}.$$656_{\pm.000}.$
Relations $72$$491_{\pm.016}.$$551_{\pm.016}.$$509_{\pm.032}.$$481_{\pm.021}.$$592_{\pm.008}.$$500_{\pm.037}.$
Overall $145$$446_{\pm.006}.$$479_{\pm.029}.$$560_{\pm.011}.$$538_{\pm.010}.$$566_{\pm.009}.$$589_{\pm.018}.$
Table 9: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
D.2 Experiments with Open-Source LLMs
Models Llama2 70b Llama2 13b Llama2 7b gemma 7b
Overall Score 0.312 0.281 0.267 0.273
Table 10: Compare the performance of Questum  method with different open-source LLMS.
Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved. This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
Script #Tokens Stage PP OP Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice Questum
Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $1.782$
Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $3.152$
Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $2.440$
Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $1.858$
Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.448$
Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.320$
Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.316$
Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $0.610$
Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.372$
Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.297$
Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $3.604$
Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $2.665$
Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $\bm{17.862}$
Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
Table 11: Compare the costs in US dollars(\$) of calling OpenAI API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. #$\#Tokens$ represent the average length of each character’s script. Costs are reported for one complete gameplay and one evaluation process for each script.
In addition to experimenting with GPT-3.5-turbo-16k 0613, we also explored Llama2 (70b, 13b, 7b) and Gemma 7b. These models were tested using the scenario “Solitary Boat Firefly", and the overall results are reported in Table 10. The findings indicate that all four models scored significantly lower than GPT-3.5, presumably due to the limitations imposed by a 4k context window. This limitation likely hindered the models’ ability to encapsulate sufficient relevant information within such a constrained window. After thorough testing, we observed that despite being evaluated on a Chinese dataset, Llama2 70b primarily resorted to English conversations due to its limited proficiency in Chinese. It responded in English even to Chinese prompts. Other models struggled even more with executing the prompts as required. This limitation greatly hindered their performance in complex gaming situations, such as MMGs. Therefore, we decided to exclusively use GPT-3.5 for future experiments, given its ability to navigate these intricate scenarios.
D.3 Ablation Studies
As shown in Table 12, we can observe that the two main modules “Search with Sensors" and “Approximate with Pruner" in our algorithm both reduced the cost and improved the performance as expected. Questum  equipped with both modules gives the best performance.
Score Gameplay Costs (\$) Gameplay Times (h)
Questum 0.407 17.861 5.740
w/o Pruner 0.396 28.731 9.932
w/o Sensor 0.379 15.057 4.457
w/o Sensor & Pruner 0.357 19.144 6.577
Table 12: Ablation study of removing the Pruner and Sensor on performance, gameplay costs, and times.
Sensor Selection
Figure 6: Compare the number of sensors, score, and cost. The chart displays the range of scores for different sensor combinations, with the I-shaped markers representing the score range and the · indicating the average score.
In practice, sensors may exhibit non-linear relationships, which necessitates careful evaluation to avoid redundancy. So we assess the unique contributions of each new sensor, ensuring that every selected sensor introduces meaningful and distinct information. For MMGs, we initialise 5 sensors with domain knowledge: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible. Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator. Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime. Evidence: This sensor evaluates whether there is evidence that directly links the mentioned character to the crime scene. Background: This sensor evaluates whether there is any history of conflict, rivalry, or enmity between the mentioned character and the rests.
As shown in Figure 6, we evaluate the performance gain provided by each sensor. The benefits diminish as additional sensors contribute less new information, while the negative impact from longer input sequences grows. Moreover, adding more sensors increases computational costs. To balance sensor effectiveness with task complexity, we selected three sensors. The final combination—Emotion, Motivation, and Opportunity Assessment—offered the best performance in our experiments. The experiment was run with different combinations of sensors. For example, we ran the experiment $\binom{5}{2}$ times to cover all possible combinations of sensors when the number of sensors is $2$.
Optimal Number of Rounds and Questum  plannings
We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7. Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions. This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of IGed from conversations tends to converge. These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases. For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
Figure 7: Comparison of agents’ behaviour across different numbers of rounds, where each agent can ask a specific number of questions (denoted as $m$).
Appendix E Agent Interaction Examples
To illustrate the effectiveness of our proposed agent, Questum, we present examples of dialogue records showcasing interactions between human players and agent players. Below are four versions in which the agent-character Inspector Howard interrogates Farmer Jack about eyewitness details concerning the murder of Merchant Green. These versions correspond to different models: Werewolf, ThinkThrice, O-CoT, and Questum.
E.1 Dialogue Versions
E.1.1 Werewolf
Inspector Howard $\rightarrow$ Farmer Jack:
“Did you notice any unusual interactions between Merchant Green and other villagers?”
Farmer Jack replies:
“No, nothing in particular.”
E.1.2 O-CoT
Inspector Howard $\rightarrow$ Farmer Jack:
“Jack, that night when you went to Merchant Green’s house, did you hear any voices or special sounds?”
Farmer Jack replies:
“I didn’t hear any voices, but I did hear the sound of a door being kicked, which was very loud and abrupt.”
E.1.3 ThinkThrice
Inspector Howard $\rightarrow$ Farmer Jack:
“That night, did you notice any unusual situations around Merchant Green’s house, like arguments or someone leaving?”
Farmer Jack replies:
“I heard some sounds, but couldn’t see who it was.”
E.1.4 Questum
Inspector Howard $\rightarrow$ Farmer Jack:
“Jack, that night you saw someone leaving Merchant Green’s house, can you recall what that person was wearing, their actions, or their general direction?”
Farmer Jack replies:
“Inspector, that night it was very dark, I could only faintly see that the person was wearing dark clothes, moving very quickly towards the back mountain path. I didn’t hear any sounds from him, but his footsteps were very hurried.”
E.2 Analysis of Models
Story Progression: Questum significantly enhances story progression by providing key clues such as the perpetrator’s dark clothing, rapid movement, and the direction towards the back mountain path. This enriches the narrative by introducing specific elements that can be further investigated. In contrast, Werewolf offers minimal advancement with a vague response denying any unusual interactions. ThinkThrice slightly improves by mentioning heard sounds but lacks actionable details, while O-CoT introduces the sound of a door being kicked but does not provide additional directions for the investigation.
Questioning Ability: Questum demonstrates superior questioning ability by targeting specific details—clothing, actions, and direction—thereby guiding Farmer Jack to provide more focused and relevant information. Werewolf poses an overly broad question that fails to elicit detailed responses. ThinkThrice narrows the focus slightly but remains too general, and O-CoT asks a more specific question but does not explore as many facets as Questum.
Answering Ability: Responses in Questum are notably detailed, offering observations about the perpetrator’s appearance and behavior, which are crucial for the investigation. This contrasts with Werewolf, where the answer is too brief and lacks specific observations. ThinkThrice provides a partial detail about hearing sounds but fails to elaborate further, while O-CoT mentions the sound of a door being kicked but does not add more context or clues.
Role Immersion: Questum excels in role immersion by portraying Farmer Jack as a tense and observant witness, recalling specific details under pressure. This adds depth to his character and makes the interaction more engaging. Werewolf’s response lacks emotional depth, presenting a mechanical denial. ThinkThrice, while slightly more descriptive, still does not convey significant emotional or psychological engagement. O-CoT includes an abrupt sound but does not enhance the character’s emotional state effectively.
Questum outperforms Werewolf, ThinkThrice, and O-CoT across all evaluated dimensions. Its precise and guided questioning prompts comprehensive and relevant answers, thereby advancing the story effectively. Additionally, the detailed responses enhance character immersion, making the narrative more compelling and believable. The comparison underscores the importance of strategic dialogue design in storytelling, where targeted interactions can significantly enrich the narrative experience.